the function for object dropped from a height where t is the time in seconds, h is the height in feet at time t, and h is the initial height
h(t)=-16t^2+h , height is 50ft I need to know how to figure out how
long it would take to fall to the ground

Answer by stanbon(75887)

h(t)=-16t^2+h , height is 50ft I need to know how to figure out how
long it would take to fall to the ground
----
When it hits the ground its height is zero.
---
Solve::
-16t^2 + 50 = 0
--------
16t^2 = 50
-------
t^2 = 50/16
----
time = 1.77 seconds
----------------
Cheers,
Stan H.
------------